Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Five people, Ada, Ben, Cathy, Dan, and Eliza, are

Expert replies
by Vincen » Tue Nov 21, 2017 8:31 am
Five people, Ada, Ben, Cathy, Dan, and Eliza, are lining up for a photo. What is the probability that none of the three girls will stand next to one another?

(A) 0.1
(B) 0.2
(C) 0.3
(D) 0.4
(E) 0.5

The OA is A .

I don't know what formula should I set. Can any expert give me some help? Please.
Join the discussion
Source: — Problem Solving |

by EconomistGMATTutor » Tue Nov 21, 2017 9:37 am
Hello Vincen.

Let's take a look at your question.

The only option that none of the three girls will stand next to one another is that they stand in the first, third and fifth places.

So, they have 3! options to stand. And the boys has 2! options. The total of possible options is 5!.

So, the answer is $$\frac{3!\cdot2!}{5!}=\frac{1}{10}=0.1.$$

The correct answer is A .

I hope this explanation may help you.

I'm available if you'd like a follow up.
GMAT Prep From The Economist
We offer 70+ point score improvement money back guarantee.
Our average student improves 98 points.

Image
Join the discussion

by [email protected] » Tue Nov 21, 2017 12:25 pm
Hi Vincen,

We're given 3 girls and 2 boys and told to put them in a line. We're asked for the probability that none of the three girls will stand next to one another. This question can be approached in a number of different ways. Here's how you can use permutations and probability to get to the correct answer:

With 5 people, there are 5! = 120 possible arrangements. Lining up the 5 people, there's just one option for the 3 girls to NOT stand next to one another:

GBGBG = (3)(2)(2)(1)(1) = 12 options

Thus, the probability of the 3 girls NOT standing next to one another is 12/120 = 1/10 = 10% = .1

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by GMATGuruNY » Tue Nov 21, 2017 1:16 pm
Vincen wrote:Five people, Ada, Ben, Cathy, Dan, and Eliza, are lining up for a photo. What is the probability that none of the three girls will stand next to one another?

(A) 0.1
(B) 0.2
(C) 0.3
(D) 0.4
(E) 0.5
For the 3 girls to be separated, they must occupy the first, third and fifth positions, as follows:
G_G_G.

P(girl occupies the first position) = 3/5. (Of the 5 children, 3 are girls.)
P(girl occupies the third position) = 2/4. (Of the 4 remaining children, 2 are girls, since 1 of the 3 girls occupies the first position.)
P(girl occupies the fifth position) = 1/3. (Of the 3 remaining children, only 1 is a girl, since 2 of the 3 girls occupy the first and third positions.)
To combine these probabilities, we multiply:
3/5 * 2/4 * 1/3 = 1/10.

The correct answer is A.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion